### v1

Abstract
Threshold-Computation-in-the-Head (TCitH) and VOLE-in-the-Head (VOLEitH), two recent developments of the MPC-in-the-Head (MPCitH) paradigm, have significantly improved the performance of digital signature schemes in this framework.
In this note, we embed the restricted decoding problem, denoted by $\mathcal{E}$-SDP, within these frameworks. We propose a structurally simple modeling that achieves competitive signature sizes. Specifically, by instantiating the restricted decoding problem with the same hardness assumption underlying CROSS, we reduce sizes by more than a factor of two compared to the NIST submission. Moreover, we observe that ternary full-weight decoding, closely related to the hardness assumption underlying WAVE, is a restricted decoding problem. Using ternary full-weight decoding, we obtain signature sizes comparable to the smallest MPCitH-based candidates in the NIST competition.
[…]
1 Introduction
Following the standardization of the signatures CRYSTALS-Dilithium, FALCON, and SPHINCS+, NIST issued a call for additional digital signature schemes in September 2022. This initiative, to which 40 proposals were submitted, has the intention of diversifying the portfolio of post-quantum signature standards. In October 2024, NIST announced that fourteen submissions are going to proceed to round two. Among these, six are based on the Threshold Computation in-the-Head (TCitH) and the VOLE in-the-head (VOLEitH) frameworks [1, 2, 3, 5, 12, 14]. These recent specializations of the Multi-Party Computation (MPC) in-the-Head (MPCitH) paradigm offer significant performance improvements compared to previous constructions.
1.0.1 MPCitH.
Originally proposed in 2007 by Ishai, Kushilevitz, Ostrovsky, and Sahai [23], the MPCitH techniques have reshaped the landscape of post-quantum digital signatures. In a nutshell, the MPCitH approach constructs digital signatures by combining secure multi-party computation (MPC) with the Fiat–Shamir transform. Emulating an MPC protocol, the prover (i.e., the signer) performs computations on shares of a secret. A zero-knowledge proof of knowledge is obtained by opening a verifier-selected subset of the parties, which is checked for consistency.
1.0.2 TCitH and VOLEitH.
The TCitH framework [21] instantiates the MPCitH paradigm with an $\ell$-out-of-$n$ threshold secret sharing. Alternatively, the TCitH framework may be described in the Polynomial Interactive Oracle Proofs (PIOP) formalism, as presented in [20]. This formalism also encompasses the VOLEitH framework, originally introduced as a technique to compile VOLE-based zero-knowledge protocols into publicly verifiable protocols, from which signatures can be obtained via the Fiat–Shamir transform [13]. Using the TCitH and VOLEitH frameworks, performant signatures can be obtained from various hardness assumptions, in particular from coding theory. Submissions to the NIST competition utilize Hamming-metric [3] and rank-metric decoding problems [2, 5], as well as the permuted kernel problem [1].
1.0.3 Our Contribution.
In this work, we consider the restricted decoding problem, called $\mathcal{E}$-SDP, underlying the CROSS signature scheme. We present a modeling of $\mathcal{E}$-SDP that allows constructing signatures in the TCitH and VOLEitH frameworks. While structurally simple, the proposed modeling achieves competitive performance. In fact, for the exact same decoding problem underlying CROSS, we reduce signature size by more than a factor of two. This shows that, by replacing the simple CVE-based proof of knowledge of CROSS with recent variants of the MPCitH framework, signature sizes can be reduced, albeit at the cost of a significantly more complex construction. Further, we consider the problem of ternary full-weight decoding, a special case of the hardness assumption underlying the WAVE signature scheme [10, 19]. Here, the proposed modeling achieves signature sizes that are comparable to the smallest ones of the NIST competition.
1.0.4 Organization.
Section 2.0.4 provides a concise review of the TCitH and VOLEitH frameworks. Table 3 presents the proposed modeling of $\mathcal{E}$-SDP, and discusses the achievable signature sizes.
1.0.5 Artifacts.
The accompanying repository [16] contains the code to reproduce the provided parameters.
2 Signatures from Threshold Computation in the Head
The VOLE-in-the-Head (VOLEitH) framework [13] and the Threshold-Computation-in-the-Head (TCitH) framework [21] are recent specializations of the MPC-in-the-Head paradigm [23]. Since we build our contribution in these frameworks, we provide a concise overview following the Polynomial Interactive Oracle Proof (PIOP) formalism presented in [20].
2.0.1 Notation.
For an integer $m\in\mathbb{Z}$, we denote $m]=\{1,\ldots,m\}$. Let $\mathbb{F}$ be the finite field with $p$ elements, where $p$ is a prime number, and $\mathbb{K}$ its degree-$\mu$ field extension. We denote by $\mathbb{F}[x_{1},\ldots,x_{n}]_{\leq d}$ the set of polynomials in variables $\mathbf{x}=(x_{1},\ldots,x_{n})$ with coefficients in $\mathbb{F}$ and total degree at most $d$. Vectors are denoted with bold lowercase letters, and matrices with bold uppercase letters. Polynomials are denoted in plain lowercase letters, and vectors of polynomials in plain uppercase letters. For a vector of polynomials $P(X)\in\mathbb{F}[X]_{\leq d}^{n}$, we denote by $P(\infty)\in\mathbb{F}^{n}$ the vector of leading (degree-$d$) coefficients.
2.0.2 TCitH in a Nutshell.
The general zero-knowledge proof of knowledge obtained from the TCitH framework is provided in Figure 1.
[…]
Figure 1: Five-pass zero-knowledge proof of knowledge.
The prover attempts to convince the verifier that they know a witness $\mathbf{w}\in\mathbb{F}^{n}$ satisfying $F\left(\mathbf{w}\right)=\mathbf{0}\in\mathbb{F}^{m}$ for a One-Way Function (OWF) $F$, described by the system of polynomial relations $F=(f_{1},\ldots,f_{m})\in\big(\mathbb{F}\left[x_{1},\ldots,x_{n}\right]_{\leq d}\big)^{m}$. To this end, the prover uses a GGM-based Vector Commitment (VC) scheme [11, 13] to sample vectors of random polynomials
$P_{u}\in\big(\mathbb{K}\left[X\right]_{\leq d-1}\big)^{\eta}\quad\text{and}\quad P_{w}\in\big(\mathbb{K}\left[X\right]_{\leq 1}\big)^{n}\quad\text{with}\quad P_{w}\left(\infty\right)\in\mathbb{F}^{n},$
which also generates a commitment $\mathsf{com}$ to $P_{w},P_{u}$ and the opening key $\mathsf{key}$. To embed the witness into $P_{w}$, the prover computes the offset $\boldsymbol{\Delta}_{\mathbf{w}}=\mathbf{w}-P_{w}\left(\infty\right)\in\mathbb{F}^{n}$, for which $\tilde{P}_{w}\left(X\right)=P_{w}\left(X\right)+\boldsymbol{\Delta}_{\mathbf{w}}\cdot X$ has leading [^1] Embedding the witness into the constant coefficient is also possible; see [14]. coefficients $\mathbf{w}$. The offset $\boldsymbol{\Delta}_{\mathbf{w}}$ and commitment $\mathsf{com}$ are then sent to the verifier. Next, the polynomial relation $F$ is evaluated at $\tilde{P}_{w}\left(X\right)$. Specifically, for $i\in[m]$ and $\ell\leq d$, denote by $f_{i,\ell}(\mathbf{x})$ the degree-$\ell$ homogeneous component of $f_{i}(\mathbf{x})$. Then, $F(\tilde{P}_{w})(X)$ is defined as
$F(\tilde{P}_{w})(X)\coloneqq\left(\sum_{\ell}X^{d-\ell}f_{1,\ell}\left(\tilde{P}_{w}\left(X\right)\right),\ldots,\sum_{\ell}X^{d-\ell}f_{m,\ell}\left(\tilde{P}_{w}\left(X\right)\right)\right).$
Then $F(\tilde{P}_{w})(\infty)=F\left(\mathbf{w}\right)$, i.e., the embedded witness satisfies the polynomial relation $F$ if and only if $F(\tilde{P}_{w})(X)$ has degree at most d1d-1. Hence, $F(\tilde{P}_{w})(X)$ consists of $m$ polynomials of degree d1d-1. The number of polynomials is reduced through batching. To this end, the verifier samples a random matrix $\boldsymbol{\Gamma}\in\mathbb{K}^{\eta\times m}$ and the prover computes $\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w})$, thereby reducing the number of polynomials from $m$ to $\eta$ while preserving the degree with overwhelming probability. The verifier masks $\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w})(X)$ by computing
$Q(X)=\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w})(X)+P_{u}\left(X\right),$
where $P_{u}\left(X\right)$ is the length-$\eta$ vector of random polynomials of degree (d1)(d-1) generated by the VC scheme. Sending the masked polynomial vector $Q(X)$ to the verifier does not reveal any information about the witness. To verify correctness, the verifier challenges the prover at a random evaluation point $r$ in the evaluation domain $\Omega\subseteq\mathbb{K}$ of size $N=\left\lvert\Omega\right\rvert$. The prover produces the corresponding opening $\mathsf{opening}_{r}$ using its opening key $\mathsf{key}$. The verifier checks $\mathsf{opening}_{r}$ against $\mathsf{com}$ and, in case of success, computes the evaluations $P_{w}\left(r\right)\in\mathbb{K}^{\eta}$ and $P_{u}\left(r\right)\in\mathbb{K}^{n}$. With these evaluations, the verifier computes $P_{u}\left(r\right)+\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w}\left(r\right))$ and checks whether it matches $Q\left(r\right)$, where $Q(X)$ is the polynomial vector received from the prover.
Remark 1
In [14], also a three-pass variant of Figure 1 is given, which is obtained by skipping the batching step. This simplifies the protocol, but increases the required communication, which is why we omit this variant.
2.0.3 TCitH signature scheme.
The proof system above has a soundness error given by
$\varepsilon=\frac{1}{p^{\mu\cdot\eta}}+\left(1-\frac{1}{p^{\mu\cdot\eta}}\right)\frac{d}{N}.$
For a target security of $\lambda$ bits, $\tau$ parallel repetitions are performed to amplify soundness below $2^{-\lambda}$. The Fiat–Shamir transform (enhanced by a $w$-bit proof of work, known as grinding) yields a signature scheme with the following properties.
Property 1
For grinding parameter $w$, let the proof system parameters satisfy
$N\leq p^{\mu},\qquad p^{\mu\cdot\eta}\geq 2^{\lambda},\quad\text{and}\quad\left(\frac{N}{d}\right)^{\tau}\geq 2^{\lambda-w}.$
Then the resulting signature scheme is EUF-CMA-secure [14]. Signatures are of the form $\sigma=(\sigma_{\mathrm{sym}},\sigma_{\mathbf{w}},\sigma_{F})$, where $\sigma_{\mathbf{w}}$ corresponds to the witness, $\sigma_{F}$ to the OWF $F$, and $\sigma_{\mathrm{sym}}$ to the VC scheme. The components’ sizes [^2] In general, the sizes depend on the specific instantiation of the VC scheme [22], and optimizations not described here can be applied. The formulas given follow [14], to which we refer the reader for further details. are given by
$\left\lvert\sigma_{\mathrm{sym}}\right\rvert$ $=\tau\cdot\lambda\cdot(\log_{2}N+1)+5\lambda+32,$
$\left\lvert\sigma_{\mathbf{w}}\right\rvert$ $=\tau\cdot n\cdot\log_{2}p,$
$\left\lvert\sigma_{F}\right\rvert$ $=\tau\cdot\eta\cdot(d-1)\cdot\mu\cdot\log_{2}p.$
2.0.4 VOLEitH.
The VOLEitH framework is closely related to the TCitH framework described above. Instead of performing $\tau$ parallel repetitions, VOLEitH employs a technique that allows sampling $r$ from an evaluation domain $\Omega$ of size
$\left\lvert\Omega\right\rvert=N^{\tau}\approx 2^{\lambda}\leq\left\lvert\mathbb{K}\right\rvert=p^{\rho}.$
Avoiding parallel repetitions reduces the soundness error, which in turn enables shorter signatures. However, this comes at the cost of a consistency check, an additional round of interaction between the prover and the verifier. For further details, we refer to [12, 13].
By parameterizing the consistency check with $B$ and the state-of-the-art VC scheme [11] with $T_{\mathrm{open}}$, a signature scheme with the following properties can be constructed.
Property 2
Let the VOLEitH proof system be parameterized such that
$\frac{N^{\tau}}{d}\geq 2^{\lambda-w}\quad\text{and}\quad p^{\rho}\geq 2^{\lambda},\quad\text{with}\quad\rho=\tau\cdot\left\lceil\log_{p}N\right\rceil,\quad B=\left\lceil\frac{16}{\log_{2}p}\right\rceil.$
Then the resulting signature scheme is EUF-CMA-secure [12], and its signatures have size [^3] The provided formulas follow the specific design choices made in [3]. $\left\lvert\sigma\right\rvert=\left\lvert\sigma_{\mathrm{sym}}\right\rvert+\left\lvert\sigma_{\mathbf{w}}\right\rvert+\left\lvert\sigma_{F}\right\rvert$, where
$\left\lvert\sigma_{\mathrm{sym}}\right\rvert$ $=\tau\cdot(\rho+B)\log_{2}p+\tau\cdot 2\lambda+T_{\mathrm{open}}\cdot\lambda+4\lambda+32,$
$\left\lvert\sigma_{\mathbf{w}}\right\rvert$ $=\tau\cdot n\cdot\log_{2}p,$
$\left\lvert\sigma_{F}\right\rvert$ $=\tau\cdot(d-1)\cdot\rho\cdot\log_{2}p.$
Similar to the TCitH framework, which requires $\eta\cdot\mu\cdot\log_{2}p\geq\lambda$, the VOLEitH framework requires $\rho\cdot\log_{2}p\geq\lambda$. However, VOLEitH allows for a smaller $\tau$, owing to its reduced soundness error, arguably at the cost of a more complex scheme.
3 Modeling Restricted Decoding Problems
In this section, we present a so-called modeling for the restricted decoding problem, i.e., we express the OWF as a polynomial relation that can be used in the TCitH and VOLEitH frameworks. Before doing so, we briefly recap some basic notions from coding theory.
3.0.1 Coding theory.
A linear code over $\mathbb{F}$ of length $n$ is a $k$-dimensional subspace of $\mathbb{F}^{n}$. The code can be represented by a parity-check matrix $\mathbf{H}\in\mathbb{F}^{(n-k)\times n}$ such that $\mathbf{c}\mathbf{H}^{\top}=\mathbf{0}$ for every codeword $\mathbf{c}$. We denote by r=nkr=n-k the redundancy of the code. The syndrome of a vector $\mathbf{v}$ (through $\mathbf{H}$) is $\mathbf{s}=\mathbf{v}\mathbf{H}^{\top}\in\mathbb{F}^{r}$. The Hamming weight (from now on, only weight) of a vector $\mathbf{v}$ is the number of non-zero entries in $\mathbf{v}$.
3.0.2 Restricted decoding assumption.
The (general) restricted Syndrome Decoding Problem ($\mathcal{E}$-SDP) [8] replaces the Hamming-weight constraint of the classical SDP with the requirement that each entry must lie in the fixed subset $\mathcal{E}$ of the underlying finite field $\mathbb{F}$. Formally, it is defined as follows:
Definition 1($\mathcal{E}$-SDP)
Let $\mathcal{E}\subset\mathbb{F}$ denote the restriction, $\mathbf{H}\in\mathbb{F}^{r\times n}$ the parity-check matrix, and $\mathbf{s}\in\mathbb{F}^{r}$ the syndrome. The $\mathcal{E}$-SDP instance $(\mathbf{H},\mathbf{s})$ is solved by $\mathbf{e}\in\mathcal{E}^{n}$ with $\mathbf{e}\mathbf{H}^{\top}=\mathbf{s}$.
We refer to the size of the restriction as $z=\left\lvert\mathcal{E}\right\rvert$. Different choices of $\mathcal{E}$ are possible, and all of them lead to an NP-complete problem [9]. Restrictions of the form $\mathcal{E}=\{-a,\ldots,a\}$ for $a\ll p$ yield a problem similar in nature to learning with errors (LWE) [8], while [9] considers cyclic subgroups of $\mathbb{F}^{*}=\mathbb{F}\setminus\{0\}$. In general, any choice of restriction requires a dedicated analysis of the problem’s concrete computational hardness. Therefore, we focus on specific choices that have undergone careful analysis.
CROSS restriction.
The signature scheme CROSS [6] uses the restriction $\mathcal{E}=\{2^{i}\mid i\in[7]\}$ within $\mathbb{F}$ of size $p=127$. We are going to refer to this setting as CROSS-SDP, the concrete hardness of which is analyzed in detail in [7] and confirmed in [15]. For instance, the parameters $n=127$ and $k=76$ provide a suitable choice for NIST security category 1.
Ternary decoding with full weight.
The signature scheme WAVE [19, 10] relies on the ternary syndrome decoding with high weight. For state-of-the-art solvers [17], parameterizing the problem such that the error is full-weight and unique, maximizes the concrete hardness (a setting in which WAVE cannot sign successfully). This corresponds to $\mathcal{E}$-SDP with $\mathcal{E}=\{1,2\}$ and $p=3$, to which we are going to refer as Ternary-SDP. For $k\approx 0.369\cdot n$, the solver presented in [17] requires $\mathcal{O}\left(2^{0.247\cdot n}\right)$ operations, up to a polynomial factor. In particular, the parameters $n=579$ and $k=213$ are shown to achieve NIST security category 1.
3.0.3 Modeling $\mathcal{E}$-SDP.
The linearity of the syndrome equations allows a standard optimization that reduces the witness size from $n$ to $k$. We assume that the parity-check matrix is in systematic form [^4] This assumption, also made e.g. in [4], does not impact the hardness of $\mathcal{E}$-SDP. $\mathbf{H}=(\mathbf{A},\mathbf{1}_{r})$, where $\mathbf{A}\in\mathbb{F}^{r\times k}$ and $\mathbf{1}_{r}\in\mathbb{F}^{r\times r}$ denotes the identity matrix. Then, $\mathbf{e}\in\mathbb{F}^{n}$ satisfies $\mathbf{e}\mathbf{H}^{\top}=\mathbf{s}$ if and only if it is of the form $\mathbf{e}=(\mathbf{w},\mathbf{s}-\mathbf{w}\mathbf{A}^{\top})$ with $\mathbf{w}\in\mathbb{F}^{k}$. That is, the witness is defined as a length-$k$ partial error vector that is expanded to a length-$n$ error vector so that the syndrome equations are satisfied.
[…]
Overall, we obtain the system of polynomial constraints $F(\mathbf{x})=(f_{1},\ldots,f_{n})\in\mathbb{F}[x_{1},\ldots,x_{k}]^{n}$ with
$f_{i}(\mathbf{x})=f_{\mathcal{E}}\left(x_{i}\right)\quad\text{for }i\in[k],\quad\text{and}\quad f_{k+i}(\mathbf{x})=f_{\mathcal{E}}\left(s_{i}-\langle\mathbf{a}_{i},\mathbf{x}\rangle\right)\quad\text{for }i\in[r],$ (1)
where $\mathbf{a}_{i}\in\mathbb{F}^{k}$ denotes the $i$-th row of $\mathbf{A}$. The correctness and degree of this modeling are stated in the following proposition.
Proposition 1
The polynomial relation $F$ given in Equation 1 provides a degree-$z$ modeling of $\mathcal{E}$-SDP: If $\mathbf{w}\in\mathbb{F}^{k}$ satisfies $F\left(\mathbf{w}\right)=\mathbf{0}$, then the error vector $\mathbf{e}=(\mathbf{w},\mathbf{s}-\mathbf{w}\mathbf{A}^{\top})$ solves the $\mathcal{E}$-SDP instance $(\mathbf{H},\mathbf{s})$ with $\mathbf{H}=(\mathbf{A},\mathbf{1}_{r})$.
Proof
For $i\in[k]$, $f_{i}(\mathbf{x})$ has degree $z$. The composition of a degree-$z$ polynomial and a linear polynomial is also of degree $z$; and, thus, $f_{k+i}(\mathbf{x})$ and $F(\mathbf{x})$ have degree $z$ as well. Let $\mathbf{w}$ satisfy $F\left(\mathbf{w}\right)=\mathbf{0}$. Then, for $\mathbf{e}=(\mathbf{w},\mathbf{s}-\mathbf{w}\mathbf{A}^{\top})$, it holds
$\mathbf{e}\mathbf{H}^{\top}=(\mathbf{w},\mathbf{s}-\mathbf{w}\mathbf{A}^{\top})(\mathbf{A},\mathbf{1}_{r})^{\top}=\mathbf{w}\mathbf{A}^{\top}+\mathbf{s}-\mathbf{w}\mathbf{A}^{\top}=\mathbf{s},$
i.e., the syndrome equations are satisfied. Further, $f_{\mathcal{E}}\left(e_{i}\right)=0$ for all entries of $\mathbf{e}$, i.e., $\mathbf{e}\in\mathcal{E}^{n}$. ∎
3.0.4 Performance in TCitH framework.
From the provided modeling of $\mathcal{E}$-SDP, a digital signature is derived using the TCitH framework [21], as summarized in Section 2.0.4. The chosen parameters and the obtained signature sizes are provided in Table 1. The parameters $\tau$, $N$, and $w$ of the proof system are selected following [14] to allow for comparable runtime performance of the VC scheme. To simplify implementation, the smallest extension degree $\mu=[\mathbb{K}:\mathbb{F}]$ satisfying the requirement $N\leq\left\lvert\mathbb{K}\right\rvert=\left\lvert\mathbb{F}\right\rvert^{\mu}$ is chosen. This, in turn, determines the batching parameter $\eta$.
Table 1: Parameters of signatures based on $\mathcal{E}$-SDP obtained via the TCitH framework. The proof system parameters are selected analogously to those in MQOM [14] to ensure comparable runtime performance of the VC scheme.
Security
category
$\mathcal{E}$-SDP Proof System Performance
[…]
NIST 1
$127$
77
$127$
$76$
$24$ $256$ 22 $10$ 44 fast $7650$
$15$ $2048$ 22 $10$ 66 short $5533$
33
22
$579$
$213$
$17$ $256$ 66 $14$ 99 fast $3533$
$12$ $2048$ 77 $12$ 88 short $3095$
NIST 3
$127$
77
$187$
$111$
$36$ $256$ 22 $14$ 66 fast 16 67516\,675
$23$ $2048$ 22 $14$ 44 short 12 35412\,354
33
22
$839$
$309$
$27$ $256$ 66 2121 33 fast $8284$
$18$ $2048$ 77 $18$ $12$ short $6860$
NIST 5
$127$
77
$251$
$150$
$48$ $256$ 22 1919 77 fast 29 83929\,839
$31$ $2048$ 22 1919 33 short 22 30522\,305
33
22
$1102$
$406$
$36$ $256$ 66 $27$ 44 fast 14 58414\,584
$25$ $2048$ 77 $24$ 66 short 12 60812\,608
For Ternary-SDP, parameterized to achieve NIST security category 1, a signature size as small as $3.1\text{\,}\mathrm{kB}$ is possible. Switching to CROSS-SDP, the signature size increases to 5.5 $5\text{\,}\mathrm{kB}.$ This is mainly due to the increased degree of the modeling and a larger witness size. Nevertheless, we obtain a significant reduction compared to CROSS [6], for which the smallest signature size is $12.4\text{\,}\mathrm{kB}$.
3.0.5 Performance in VOLEitH framework.
Table 2 provides an overview of the performances that can be achieved if TCitH is replaced by the VOLEitH framework. Parameters of the proof system are chosen analogously to SDitH [3]. For Ternary-SDP, only a minor improvement in signature size is achievable. A similar effect is observed in [14]: for OWFs with small degree and small witness size, the structurally simpler TCitH framework remains competitive. On the other hand, for CROSS-SDP, which has a degree of $d=z=7$, VOLEitH enables a reduction in signature size of about $1\text{\,}\mathrm{kB}$ for NIST security category 1, and of up to $6\text{\,}\mathrm{kB}$ for NIST security category 5.
Table 2: Parameters of signatures based on $\mathcal{E}$-SDP obtained via the VOLEitH framework. The proof system parameters are selected analogously to those in SDitH [3] to ensure comparable runtime performance of the VC scheme.
Security
category
$\mathcal{E}$-SDP Proof System Performance
[…]
NIST 1
$127$
77
$127$
$76$
$16$ $256$ $32$ $101$ 33 fast $6432$
$11$ $2048$ $22$ $107$ $10$ short $4372$
33
22
$579$
$213$
$16$ $256$ $96$ $101$ 11 fast $3515$
$11$ $2048$ $81$ $107$ 88 short $2974$
NIST 3
$127$
77
$187$
$111$
$24$ $256$ $48$ $153$ 33 fast 14 35914\,359
$16$ $4096$ $32$ $157$ 33 short $9361$
33
22
$839$
$309$
$24$ $256$ $144$ $153$ 11 fast $7816$
$16$ $4096$ $128$ $157$ 11 short $6463$
NIST 5
$127$
77
$251$
$150$
$32$ $256$ $64$ $206$ 33 fast 25 57325\,573
2121 $4096$ $42$ $216$ 77 short 16 58916\,589
33
22
$1102$
$406$
$32$ $256$ $192$ $206$ 11 fast 13 85113\,851
2121 $4096$ $168$ $216$ 55 short 11 52111\,521
3.0.6 Comparison with NIST submissions.
A comparison of the achieved signature sizes with submissions to NIST’s call for additional digital signature schemes [24] is provided in Table 3. As can be seen, the TCitH and VOLEitH frameworks can be combined with a wide range of OWFs. The proposed modeling of $\mathcal{E}$-SDP is competitive compared to the state-of-the-art: for the Ternary-SDP, we obtain a signature size which is very close to those of the smallest proposals that utilize the TCitH or the VOLEitH framework with an alternative hardness assumption. The sizes obtained from CROSS-SDP are slightly larger than those of SDitH and PERK, which also use modelings of degree greater than two, but smaller than $z=7$. The code-based proposals CROSS and LESS target other trade-offs: CROSS uses a simple CVE-like protocol [18, 6], and LESS achieves a small signature size by using multiple public keys (resulting in a public-key size that is multiple times larger than those of the other schemes shown).
Table 3: Comparison of the obtained signature sizes with NIST submissions. The shown sizes correspond to the short optimization tradeoff; runtime may vary.
Scheme Hardness Assumption Design Principle $\left\lvert\sigma\right\rvert$ in $\mathrm{kB}$]
this work CROSS-SDP VOLEitH $4.4$
Ternary-SDP TCitH $3.1$
CROSS CROSS-$\mathcal{E}$-SDP CVE $12.4$
LESS Code Equivalence sigma protocol $1.3$
SDitH Syndrome Decoding VOLEitH $3.7$
RYDE Rank Decoding TCitH $3.1$
Mirath MinRank TCitH $3.0$
PERK Permuted Kernel VOLEitH $3.5$
MQOM Multivariate Quadratic TCitH $2.8$
FEAST AES VOLEitH $4.5$
AES-EM $3.9$
4 Conclusion
This paper presents a polynomial modeling of the restricted decoding problem $\mathcal{E}$-SDP, enabling the construction of digital signatures in the TCitH and the VOLEitH frameworks. When instantiated with ternary full-weight decoding, we obtain signature sizes that are competitive with the state-of-the-art NIST candidates. For the restriction used by the CROSS signature scheme, the sizes are significantly reduced compared to CROSS.

### v2

@@
-
+TCitH- and VOLEitH-based Signatures
 Abstract
-Threshold-Computation-in-the-Head (TCitH) and VOLE-in-the-Head (VOLEitH), two recent developments of the MPC-in-the-Head (MPCitH) paradigm, have significantly improved the performance of digital signature schemes in this framework.
-In this note, we embed the restricted decoding problem, denoted by $\mathcal{E}$-SDP, within these frameworks.
-We propose a structurally simple modeling that achieves competitive signature sizes.
+Threshold-Computation-in-the-Head (TCitH) and VOLE-in-the-Head (VOLEitH), two recent developments of the MPC-in-the-Head (MPCitH) paradigm, have significantly improved the performance of digital signature schemes.
+This work embeds the restricted decoding problem within these frameworks: we propose a structurally simple modeling that achieves competitive signature sizes.
 Specifically, by instantiating the restricted decoding problem with the same hardness assumption underlying CROSS, we reduce sizes by more than a factor of two compared to the NIST submission.
@@
 Using ternary full-weight decoding, we obtain signature sizes comparable to the smallest MPCitH-based candidates in the NIST competition.
-[…]
-1 Introduction
-Following the standardization of the signatures CRYSTALS-Dilithium, FALCON, and SPHINCS+, NIST issued a call for additional digital signature schemes in September 2022.
-This initiative, to which 40 proposals were submitted, has the intention of diversifying the portfolio of post-quantum signature standards.
-In October 2024, NIST announced that fourteen submissions are going to proceed to round two.
-Among these, six are based on the Threshold Computation in-the-Head (TCitH) and the VOLE in-the-head (VOLEitH) frameworks [1, 2, 3, 5, 12, 14].
-These recent specializations of the Multi-Party Computation (MPC) in-the-Head (MPCitH) paradigm offer significant performance improvements compared to previous constructions.
-1.0.1 MPCitH.
+I Introduction
+Post-quantum signatures.
+Following the standardization of the signatures Dilithium (ML-DSA), FALCON (FN-DSA), and SPHINCS+ (SLH-DSA), NIST issued a call for additional digital signature schemes in September 2022.
+In October 2024, NIST announced that fourteen submissions proceeded to the second round.
+Among these, six are based on the Threshold Computation in-the-Head (TCitH) and the Vector Oblivious Linear Evaluation (VOLE) in-the-head (VOLEitH) technique [1, 2, 3, 4, 5, 6].
+These recent specializations of the Multi-Party Computation (MPC) in-the-Head (MPCitH) paradigm offer significant performance improvements.
+MPCitH.
 Originally proposed in 2007 by Ishai, Kushilevitz, Ostrovsky, and Sahai [7], the MPCitH techniques have reshaped the landscape of post-quantum digital signatures.
-In a nutshell, the MPCitH approach constructs digital signatures by combining secure multi-party computation (MPC) with the Fiat–Shamir transform.
+In a nutshell, the MPCitH approach constructs digital signatures by combining secure MPC with the Fiat–Shamir transform.
 Emulating an MPC protocol, the prover (i.e., the signer) performs computations on shares of a secret.
 A zero-knowledge proof of knowledge is obtained by opening a verifier-selected subset of the parties, which is checked for consistency.
-1.0.2 TCitH and VOLEitH.
+TCitH and VOLEitH.
 The TCitH framework [8] instantiates the MPCitH paradigm with an $\ell$-out-of-$n$ threshold secret sharing.
@@
 This formalism also encompasses the VOLEitH framework, originally introduced as a technique to compile VOLE-based zero-knowledge protocols into publicly verifiable protocols, from which signatures can be obtained via the Fiat–Shamir transform [10].
-Using the TCitH and VOLEitH frameworks, performant signatures can be obtained from various hardness assumptions, in particular from coding theory.
-Submissions to the NIST competition utilize Hamming-metric [3] and rank-metric decoding problems [2, 5], as well as the permuted kernel problem [1].
-1.0.3 Our Contribution.
-In this work, we consider the restricted decoding problem, called $\mathcal{E}$-SDP, underlying the CROSS signature scheme.
+Code-based Cryptography refers to a class of cryptographic protocol whose security relies on hard problems in coding theory, such as syndrome decoding or code equivalence.
+In the context of the ongoing NIST competition, six out of the fourteen submissions are based on different coding-theory related assumption, such as code equivalence [11], restricted decoding [12], Hamming-metric decoding [3], rank-metric decoding decoding [2, 4], and the permuted kernel problem [1].
+Among these, the last four are based on either TCitH or VOLEitH.
+Our Contribution.
+In this work, we consider the restricted decoding problem, called $\mathcal{E}$-SDP, underlying the CROSS signature scheme [12].
 We present a modeling of $\mathcal{E}$-SDP that allows constructing signatures in the TCitH and VOLEitH frameworks.
-While structurally simple, the proposed modeling achieves competitive performance.
-In fact, for the exact same decoding problem underlying CROSS, we reduce signature size by more than a factor of two.
-This shows that, by replacing the simple CVE-based proof of knowledge of CROSS with recent variants of the MPCitH framework, signature sizes can be reduced, albeit at the cost of a significantly more complex construction.
+By doing so, we reduce CROSS signature sizes by more than a factor of two, at the cost of a more complex construction.
 Further, we consider the problem of ternary full-weight decoding, a special case of the hardness assumption underlying the WAVE signature scheme [13, 14].
-Here, the proposed modeling achieves signature sizes that are comparable to the smallest ones of the NIST competition.
-1.0.4 Organization.
-Section 2.0.4 provides a concise review of the TCitH and VOLEitH frameworks. Table 3 presents the proposed modeling of $\mathcal{E}$-SDP, and discusses the achievable signature sizes.
-1.0.5 Artifacts.
-The accompanying repository [16] contains the code to reproduce the provided parameters.
-2 Signatures from Threshold Computation in the Head
-The VOLE-in-the-Head (VOLEitH) framework [13] and the Threshold-Computation-in-the-Head (TCitH) framework [21] are recent specializations of the MPC-in-the-Head paradigm [23].
-Since we build our contribution in these frameworks, we provide a concise overview following the Polynomial Interactive Oracle Proof (PIOP) formalism presented in [20].
-2.0.1 Notation.
-For an integer $m\in\mathbb{Z}$, we denote $m]=\{1,\ldots,m\}$.
-Let $\mathbb{F}$ be the finite field with $p$ elements, where $p$ is a prime number, and $\mathbb{K}$ its degree-$\mu$ field extension.
-We denote by $\mathbb{F}[x_{1},\ldots,x_{n}]_{\leq d}$ the set of polynomials in variables $\mathbf{x}=(x_{1},\ldots,x_{n})$ with coefficients in $\mathbb{F}$ and total degree at most $d$.
+Here, the proposed modeling achieves signature sizes comparable to the smallest [^1] Except for SQISign [15], which is considerably shorter and slower.
+Zero-Knowledge-based (ZK-based) schemes of the NIST competition.
+Organization.
+Section˜II provides a review of the TCitH and VOLEitH frameworks.
+Section˜III presents the proposed modeling of $\mathcal{E}$-SDP, and Section˜IV provides a detailed performance evaluation.
+The repository [16] contains code to reproduce the provided figures and parameters.
+II TCitH and VOLEitH frameworks
+[…]
+Figure 1: Five-pass zero-knowledge proof of knowledge.
+We provide a concise overview following the PIOP formalism presented in [9] used to build TCitH and VOLEitH signatures.
+Notation.
+$\lambda$ denotes the security parameter; NIST (security) categories 1, 3, and 5 correspond to $\lambda=143$, $\lambda=207$, and $\lambda=272$, respectively.
+For an integer $m$, we denote by $m]=\{1,\ldots,m\}$.
+Let $\mathbb{F}_{p}$ be the finite field with $p$ elements, where $p$ is a prime number, and $\mathbb{K}$ its degree-$\mu$ field extension.
+We denote by $\mathbb{F}_{p}[x_{1},\ldots,x_{k}]_{\leq d}$ the set of polynomials in variables $\mathbf{x}=(x_{1},\ldots,x_{k})$ with coefficients in $\mathbb{F}_{p}$ and total degree at most $d$.
 Vectors are denoted with bold lowercase letters, and matrices with bold uppercase letters.
 Polynomials are denoted in plain lowercase letters, and vectors of polynomials in plain uppercase letters.
-For a vector of polynomials $P(X)\in\mathbb{F}[X]_{\leq d}^{n}$, we denote by $P(\infty)\in\mathbb{F}^{n}$ the vector of leading (degree-$d$) coefficients.
-2.0.2 TCitH in a Nutshell.
-The general zero-knowledge proof of knowledge obtained from the TCitH framework is provided in Figure 1.
-[…]
-Figure 1: Five-pass zero-knowledge proof of knowledge.
-The prover attempts to convince the verifier that they know a witness $\mathbf{w}\in\mathbb{F}^{n}$ satisfying $F\left(\mathbf{w}\right)=\mathbf{0}\in\mathbb{F}^{m}$ for a One-Way Function (OWF) $F$, described by the system of polynomial relations $F=(f_{1},\ldots,f_{m})\in\big(\mathbb{F}\left[x_{1},\ldots,x_{n}\right]_{\leq d}\big)^{m}$.
+For a vector of polynomials in the variable $X$ $P(X)\in\mathbb{F}_{p}[X]_{\leq d}^{k}$, we denote by $P(\infty)\in\mathbb{F}_{p}^{k}$ the vector of leading (degree-$d$) coefficients.
+We write $x\propto y$ to denote that $x$ is proportional to $y$.
+TCitH.
+The general zero-knowledge proof of knowledge obtained from the TCitH framework is provided in Figure˜1.
+We give a brief overview below; for further details, see [8, 6].
+The prover attempts to convince the verifier that they know a witness $\mathbf{w}\in\mathbb{F}_{p}^{k}$ satisfying $F\left(\mathbf{w}\right)=\mathbf{0}\in\mathbb{F}_{p}^{n}$ for a One-Way Function (OWF) $F$, described by the system of polynomial relations $F=(f_{1},\ldots,f_{m})\in\big(\mathbb{F}_{p}\left[x_{1},\ldots,x_{k}\right]_{\leq d}\big)^{m}$.
 To this end, the prover uses a GGM-based Vector Commitment (VC) scheme [17, 10] to sample vectors of random polynomials
-$P_{u}\in\big(\mathbb{K}\left[X\right]_{\leq d-1}\big)^{\eta}\quad\text{and}\quad P_{w}\in\big(\mathbb{K}\left[X\right]_{\leq 1}\big)^{n}\quad\text{with}\quad P_{w}\left(\infty\right)\in\mathbb{F}^{n},$
-which also generates a commitment $\mathsf{com}$ to $P_{w},P_{u}$ and the opening key $\mathsf{key}$.
-To embed the witness into $P_{w}$, the prover computes the offset $\boldsymbol{\Delta}_{\mathbf{w}}=\mathbf{w}-P_{w}\left(\infty\right)\in\mathbb{F}^{n}$, for which $\tilde{P}_{w}\left(X\right)=P_{w}\left(X\right)+\boldsymbol{\Delta}_{\mathbf{w}}\cdot X$ has leading [^1] Embedding the witness into the constant coefficient is also possible; see [14].
+$P_{u}\in\big(\mathbb{K}\left[X\right]_{\leq d-1}\big)^{\eta}\quad\text{and}\quad P_{w}\in\big(\mathbb{K}\left[X\right]_{\leq 1}\big)^{k}$
+with $P_{w}\left(\infty\right)\in\mathbb{F}_{p}^{k}$.
+The VC also generates a commitment $\mathsf{com}$ to $P_{w},P_{u}$ and the opening key $\mathsf{key}$.
+To embed the witness $\mathbf{w}$ into $P_{w}$, the prover computes the offset $\boldsymbol{\Delta}_{\mathbf{w}}=\mathbf{w}-P_{w}\left(\infty\right)\in\mathbb{F}_{p}^{k}$, for which $\tilde{P}_{w}\left(X\right)=P_{w}\left(X\right)+\boldsymbol{\Delta}_{\mathbf{w}}\cdot X$ has leading [^2] The witness can also be embedded into the constant coefficient; see [6].
 coefficients $\mathbf{w}$.
@@
 Next, the polynomial relation $F$ is evaluated at $\tilde{P}_{w}\left(X\right)$.
-Specifically, for $i\in[m]$ and $\ell\leq d$, denote by $f_{i,\ell}(\mathbf{x})$ the degree-$\ell$ homogeneous component of $f_{i}(\mathbf{x})$.
+Specifically, for $i\in[n]$ and $\ell\leq d$, denote by $f_{i,\ell}(\mathbf{x})$ the degree-$\ell$ homogeneous component (i.e.
+the sum of all the monomials of degree $\ell$) of $f_{i}(\mathbf{x})$ .
 Then, $F(\tilde{P}_{w})(X)$ is defined as
-$F(\tilde{P}_{w})(X)\coloneqq\left(\sum_{\ell}X^{d-\ell}f_{1,\ell}\left(\tilde{P}_{w}\left(X\right)\right),\ldots,\sum_{\ell}X^{d-\ell}f_{m,\ell}\left(\tilde{P}_{w}\left(X\right)\right)\right).$
-Then $F(\tilde{P}_{w})(\infty)=F\left(\mathbf{w}\right)$, i.e., the embedded witness satisfies the polynomial relation $F$ if and only if $F(\tilde{P}_{w})(X)$ has degree at most d1d-1.
-Hence, $F(\tilde{P}_{w})(X)$ consists of $m$ polynomials of degree d1d-1.
+$\Big(\sum_{\ell}X^{d-\ell}f_{1,\ell}(\tilde{P}_{w}\left(X\right)),\ldots,\sum_{\ell}X^{d-\ell}f_{n,\ell}(\tilde{P}_{w}\left(X\right))\Big).$
+Thus, $F(\tilde{P}_{w})(\infty)=F\left(\mathbf{w}\right)$, i.e., the embedded witness satisfies the polynomial relation $F$ if and only if $F(\tilde{P}_{w})(X)$ has degree at most d1d-1.
+Hence, $F(\tilde{P}_{w})(X)$ consists of $n$ polynomials of degree d1d-1.
 The number of polynomials is reduced through batching.
-To this end, the verifier samples a random matrix $\boldsymbol{\Gamma}\in\mathbb{K}^{\eta\times m}$ and the prover computes $\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w})$, thereby reducing the number of polynomials from $m$ to $\eta$ while preserving the degree with overwhelming probability.
-The verifier masks $\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w})(X)$ by computing
-$Q(X)=\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w})(X)+P_{u}\left(X\right),$
-where $P_{u}\left(X\right)$ is the length-$\eta$ vector of random polynomials of degree (d1)(d-1) generated by the VC scheme.
+To this end, the verifier samples a random matrix $\boldsymbol{\Gamma}\in\mathbb{K}^{\eta\times n}$ and the prover computes $\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w})$, thereby reducing the number of polynomials from $n$ to $\eta$ while preserving the degree with overwhelming probability.
+The prover masks $\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w})(X)$ by computing
+$Q(X)=P_{u}\left(X\right)+\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w})(X),\vskip-2.84526pt$
+where $P_{u}\left(X\right)$ is the length-$\eta$ vector of random polynomials of degree d1d-1 generated by the VC scheme.
 Sending the masked polynomial vector $Q(X)$ to the verifier does not reveal any information about the witness.
@@
 The prover produces the corresponding opening $\mathsf{opening}_{r}$ using its opening key $\mathsf{key}$.
-The verifier checks $\mathsf{opening}_{r}$ against $\mathsf{com}$ and, in case of success, computes the evaluations $P_{w}\left(r\right)\in\mathbb{K}^{\eta}$ and $P_{u}\left(r\right)\in\mathbb{K}^{n}$.
-With these evaluations, the verifier computes $P_{u}\left(r\right)+\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w}\left(r\right))$ and checks whether it matches $Q\left(r\right)$, where $Q(X)$ is the polynomial vector received from the prover.
-Remark 1
-In [14], also a three-pass variant of Figure 1 is given, which is obtained by skipping the batching step.
+The verifier checks $\mathsf{opening}_{r}$ against $\mathsf{com}$ and, in case of success, computes the evaluations $P_{w}\left(r\right)\in\mathbb{K}^{\eta}$ and $P_{u}\left(r\right)\in\mathbb{K}^{k}$.
+Then, the verifier computes $P_{u}\left(r\right)+\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w}\left(r\right))$ and checks whether it matches $Q\left(r\right)$, where $Q(X)$ is the polynomial vector received from the prover.
+Remark 1.
+In [6], also a three-pass variant of Figure˜1 is given, which is obtained by skipping the batching step.
 This simplifies the protocol, but increases the required communication, which is why we omit this variant.
-2.0.3 TCitH signature scheme.
-The proof system above has a soundness error given by
+TCitH signature scheme. The proof system above has a soundness error given by
 $\varepsilon=\frac{1}{p^{\mu\cdot\eta}}+\left(1-\frac{1}{p^{\mu\cdot\eta}}\right)\frac{d}{N}.$
@@
 The Fiat–Shamir transform (enhanced by a $w$-bit proof of work, known as grinding) yields a signature scheme with the following properties.
-Property 1
-For grinding parameter $w$, let the proof system parameters satisfy
+Property 1.
+For grinding parameter $w$, let the TCitH proof system parameters satisfy
 $N\leq p^{\mu},\qquad p^{\mu\cdot\eta}\geq 2^{\lambda},\quad\text{and}\quad\left(\frac{N}{d}\right)^{\tau}\geq 2^{\lambda-w}.$
@@
 are given by
-$\left\lvert\sigma_{\mathrm{sym}}\right\rvert$ $=\tau\cdot\lambda\cdot(\log_{2}N+1)+5\lambda+32,$
-$\left\lvert\sigma_{\mathbf{w}}\right\rvert$ $=\tau\cdot n\cdot\log_{2}p,$
-$\left\lvert\sigma_{F}\right\rvert$ $=\tau\cdot\eta\cdot(d-1)\cdot\mu\cdot\log_{2}p.$
-2.0.4 VOLEitH.
-The VOLEitH framework is closely related to the TCitH framework described above.
-Instead of performing $\tau$ parallel repetitions, VOLEitH employs a technique that allows sampling $r$ from an evaluation domain $\Omega$ of size
+$\left\lvert\sigma_{\mathrm{sym}}\right\rvert=\tau\cdot\lambda\cdot(\log_{2}(N)+1)+5\lambda+32,$
+$\left\lvert\sigma_{\mathbf{w}}\right\rvert=\tau\cdot k\cdot\log_{2}(p),\text{ }\left\lvert\sigma_{F}\right\rvert=\tau\cdot\eta\cdot(d-1)\cdot\mu\cdot\log_{2}(p).\\$
+VOLEitH. Instead of performing $\tau$ parallel repetitions, VOLEitH employs a technique that allows sampling $r$ from an evaluation domain $\Omega$ of size
 $\left\lvert\Omega\right\rvert=N^{\tau}\approx 2^{\lambda}\leq\left\lvert\mathbb{K}\right\rvert=p^{\rho}.$
 Avoiding parallel repetitions reduces the soundness error, which in turn enables shorter signatures.
-However, this comes at the cost of a consistency check, an additional round of interaction between the prover and the verifier.
+However, this comes at the cost of an additional round of interaction between the prover and the verifier to do a a consistency check.
 For further details, we refer to [5, 10].
 By parameterizing the consistency check with $B$ and the state-of-the-art VC scheme [17] with $T_{\mathrm{open}}$, a signature scheme with the following properties can be constructed.
-Property 2
+Property 2.
 Let the VOLEitH proof system be parameterized such that
-$\frac{N^{\tau}}{d}\geq 2^{\lambda-w}\quad\text{and}\quad p^{\rho}\geq 2^{\lambda},\quad\text{with}\quad\rho=\tau\cdot\left\lceil\log_{p}N\right\rceil,\quad B=\left\lceil\frac{16}{\log_{2}p}\right\rceil.$
+$\frac{N^{\tau}}{d}\geq 2^{\lambda-w}\quad\text{and}\quad p^{\rho}\geq 2^{\lambda},$
+$\text{with}\quad\rho=\tau\cdot\left\lceil\log_{p}(N)\right\rceil,\quad B=\left\lceil\frac{16}{\log_{2}(p)}\right\rceil.$
 Then the resulting signature scheme is EUF-CMA-secure [5], and its signatures have size [^4] The provided formulas follow the specific design choices made in [3].
 $\left\lvert\sigma\right\rvert=\left\lvert\sigma_{\mathrm{sym}}\right\rvert+\left\lvert\sigma_{\mathbf{w}}\right\rvert+\left\lvert\sigma_{F}\right\rvert$, where
-$\left\lvert\sigma_{\mathrm{sym}}\right\rvert$ $=\tau\cdot(\rho+B)\log_{2}p+\tau\cdot 2\lambda+T_{\mathrm{open}}\cdot\lambda+4\lambda+32,$
-$\left\lvert\sigma_{\mathbf{w}}\right\rvert$ $=\tau\cdot n\cdot\log_{2}p,$
-$\left\lvert\sigma_{F}\right\rvert$ $=\tau\cdot(d-1)\cdot\rho\cdot\log_{2}p.$
-Similar to the TCitH framework, which requires $\eta\cdot\mu\cdot\log_{2}p\geq\lambda$, the VOLEitH framework requires $\rho\cdot\log_{2}p\geq\lambda$.
+$\left\lvert\sigma_{\mathrm{sym}}\right\rvert$ $=\tau\cdot(\rho+B)\log_{2}(p)+\tau\cdot 2\lambda+T_{\mathrm{open}}\cdot\lambda+4\lambda+32,$
+$\left\lvert\sigma_{\mathbf{w}}\right\rvert$ $=\tau\cdot k\cdot\log_{2}(p),$
+$\left\lvert\sigma_{F}\right\rvert$ $=\tau\cdot(d-1)\cdot\rho\cdot\log_{2}(p).$
+Similar to the TCitH framework, which requires $\eta\cdot\mu\cdot\log_{2}(p)\geq\lambda$, the VOLEitH framework requires $\rho\cdot\log_{2}(p)\geq\lambda$.
 However, VOLEitH allows for a smaller $\tau$, owing to its reduced soundness error, arguably at the cost of a more complex scheme.
-3 Modeling Restricted Decoding Problems
-In this section, we present a so-called modeling for the restricted decoding problem, i.e., we express the OWF as a polynomial relation that can be used in the TCitH and VOLEitH frameworks.
-Before doing so, we briefly recap some basic notions from coding theory.
-3.0.1 Coding theory.
-A linear code over $\mathbb{F}$ of length $n$ is a $k$-dimensional subspace of $\mathbb{F}^{n}$.
-The code can be represented by a parity-check matrix $\mathbf{H}\in\mathbb{F}^{(n-k)\times n}$ such that $\mathbf{c}\mathbf{H}^{\top}=\mathbf{0}$ for every codeword $\mathbf{c}$.
+TABLE I: Parameters of signatures based on $\mathcal{E}$-SDP obtained via the TCitH framework.
+The proof system parameters are selected analogously to those in MQOM [6] to ensure comparable runtime performance of the VC scheme.
+Security category $\mathcal{E}$-SDP Proof System Performance
+[…]
+NIST 1 $127$ 77 $127$ $76$ $24$ $256$ 22 $10$ 44 fast $7650$
+$15$ $2048$ 22 $10$ 66 short $5533$
+33 22 $579$ $213$ $17$ $256$ 66 $14$ 99 fast $3533$
+$12$ $2048$ 77 $12$ 88 short $3095$
+NIST 3 $127$ 77 $187$ $111$ $36$ $256$ 22 $14$ 66 fast 16 67516\,675
+$23$ $2048$ 22 $14$ 44 short 12 35412\,354
+33 22 $839$ $309$ $27$ $256$ 66 2121 33 fast $8284$
+$18$ $2048$ 77 $18$ $12$ short $6860$
+NIST 5 $127$ 77 $251$ $150$ $48$ $256$ 22 1919 77 fast 29 83929\,839
+$31$ $2048$ 22 1919 33 short 22 30522\,305
+33 22 $1102$ $406$ $36$ $256$ 66 $27$ 44 fast 14 58414\,584
+$25$ $2048$ 77 $24$ 66 short 12 60812\,608
+III Modeling Restricted Decoding Problems
+We now present a polynomial modeling for $\mathcal{E}$-SDP that can be utilized in the TCitH and VOLEitH frameworks. First, we recap some basic notions from coding theory.
+Coding theory.
+A linear code over $\mathbb{F}_{p}$ of length $n$ is a $k$-dimensional subspace of $\mathbb{F}_{p}^{n}$.
+The code can be represented by a parity-check matrix $\mathbf{H}\in\mathbb{F}_{p}^{(n-k)\times n}$ such that $\mathbf{c}\mathbf{H}^{\top}=\mathbf{0}$ for every codeword $\mathbf{c}$.
 We denote by r=nkr=n-k the redundancy of the code.
-The syndrome of a vector $\mathbf{v}$ (through $\mathbf{H}$) is $\mathbf{s}=\mathbf{v}\mathbf{H}^{\top}\in\mathbb{F}^{r}$.
+The syndrome of a vector $\mathbf{v}$ (through $\mathbf{H}$) is $\mathbf{s}=\mathbf{v}\mathbf{H}^{\top}\in\mathbb{F}_{p}^{r}$.
 The Hamming weight (from now on, only weight) of a vector $\mathbf{v}$ is the number of non-zero entries in $\mathbf{v}$.
-3.0.2 Restricted decoding assumption.
-The (general) restricted Syndrome Decoding Problem ($\mathcal{E}$-SDP) [8] replaces the Hamming-weight constraint of the classical SDP with the requirement that each entry must lie in the fixed subset $\mathcal{E}$ of the underlying finite field $\mathbb{F}$.
-Formally, it is defined as follows:
-Definition 1($\mathcal{E}$-SDP)
-Let $\mathcal{E}\subset\mathbb{F}$ denote the restriction, $\mathbf{H}\in\mathbb{F}^{r\times n}$ the parity-check matrix, and $\mathbf{s}\in\mathbb{F}^{r}$ the syndrome.
-The $\mathcal{E}$-SDP instance $(\mathbf{H},\mathbf{s})$ is solved by $\mathbf{e}\in\mathcal{E}^{n}$ with $\mathbf{e}\mathbf{H}^{\top}=\mathbf{s}$.
+Restricted decoding assumption.
+The (general) restricted Syndrome Decoding Problem ($\mathcal{E}$-SDP) [19] replaces the Hamming-weight constraint of the usual SDP with the requirement that each entry must lie in the fixed subset $\mathcal{E}$ of the underlying finite field $\mathbb{F}_{p}$.
+It is defined as follows:
+Definition 2 ($\mathcal{E}$-SDP).
+Let the restriction $\mathcal{E}\subset\mathbb{F}_{p}$, the parity-check matrix $\mathbf{H}\in\mathbb{F}_{p}^{r\times n}$, and the syndrome $\mathbf{s}\in\mathbb{F}_{p}^{r}$ be given.
+Solving $\mathcal{E}$-SDP requires to find $\mathbf{e}\in\mathcal{E}^{n}$ such that $\mathbf{e}\mathbf{H}^{\top}=\mathbf{s}$.
 We refer to the size of the restriction as $z=\left\lvert\mathcal{E}\right\rvert$.
 Different choices of $\mathcal{E}$ are possible, and all of them lead to an NP-complete problem [20].
-Restrictions of the form $\mathcal{E}=\{-a,\ldots,a\}$ for $a\ll p$ yield a problem similar in nature to learning with errors (LWE) [8], while [9] considers cyclic subgroups of $\mathbb{F}^{*}=\mathbb{F}\setminus\{0\}$.
-In general, any choice of restriction requires a dedicated analysis of the problem’s concrete computational hardness.
-Therefore, we focus on specific choices that have undergone careful analysis.
-CROSS restriction.
-The signature scheme CROSS [6] uses the restriction $\mathcal{E}=\{2^{i}\mid i\in[7]\}$ within $\mathbb{F}$ of size $p=127$.
+In this work, we mainly focus on two specific choices that have undergone careful analysis.
+CROSS restriction
+The signature scheme CROSS [12] uses the restriction $\mathcal{E}=\{2^{i}\mid i\in[7]\}$ within $\mathbb{F}_{p}$ of size $p=127$.
 We are going to refer to this setting as CROSS-SDP, the concrete hardness of which is analyzed in detail in [21] and confirmed in [22].
 For instance, the parameters $n=127$ and $k=76$ provide a suitable choice for NIST security category 1.
-Ternary decoding with full weight.
+Ternary decoding with full weight
 The signature scheme WAVE [14, 13] relies on the ternary syndrome decoding with high weight.
@@
 This corresponds to $\mathcal{E}$-SDP with $\mathcal{E}=\{1,2\}$ and $p=3$, to which we are going to refer as Ternary-SDP.
-For $k\approx 0.369\cdot n$, the solver presented in [17] requires $\mathcal{O}\left(2^{0.247\cdot n}\right)$ operations, up to a polynomial factor.
+When polynomial factors are conservatively ignored, the solver presented in [23] requires $2^{0.247\cdot n}$ operations for $k\approx 0.369\cdot n$, which implies uniqueness of the solution.
 In particular, the parameters $n=579$ and $k=213$ are shown to achieve NIST security category 1.
-3.0.3 Modeling $\mathcal{E}$-SDP.
+General $\mathcal{E}$-SDP
+In general, any choice of restriction requires a careful analysis of the problem’s concrete computational hardness.
+This makes suggesting parameters difficult if one deviates from the settings described above.
+Nevertheless, we are interested in assessing whether other parameterizations can yield an improved performance in combination with the proposed modeling.
+An upper bound on the cost of solving a general $\mathcal{E}$-SDP instance is provided by Stern’s algorithm, as adapted in [20].
+The following proposition provides a tight closed-form approximation for the cost of this solver.
+Proposition 1.
+Let $Z=\log_{2}\left(z\right)$, $P=\log_{2}(p)$, and k/n=1Z/Pk/n=1-Z/P. Then, conservatively ignoring polynomial factors and memory cost, Stern’s solver has time complexity $C$ with
+$\log_{2}\left(C\right)\approx\frac{P-Z}{2\cdot P-Z}\cdot Z\cdot n.$ (1)
+Proof.
+For k/n=1Z/Pk/n=1-Z/P, the solution $\mathbf{e}$ is unique with good probability [21]. Following [20], the cost of Stern’s solver is
+$\min_{0\leq\ell\leq r}\left\{z^{\frac{k+\ell}{2}}+\frac{z^{k+\ell}}{p^{\ell}}\right\},$
+when polynomial factors and memory cost are neglected.
+The optimal $\ell$ balances $z^{\frac{k+\ell}{2}}=\frac{z^{k+\ell}}{p^{\ell}}$; this is achieved for $k/\ell=2P/Z-1$.
+Equation˜1 follows from $C\approx z^{\frac{k+\ell}{2}}=p^{\ell}$.
+∎
+Progress in cryptanalysis of $\mathcal{E}$-SDP would require increased $n$.
+As in the case of [24], since the signature size is typically dominated by $\sigma_{\mathrm{sym}}$ for the modeling we propose, this would have only a minor impact on performance in the TCitH and VOLEitH frameworks.
+Modeling $\mathcal{E}$-SDP.
 The linearity of the syndrome equations allows a standard optimization that reduces the witness size from $n$ to $k$.
-We assume that the parity-check matrix is in systematic form [^4] This assumption, also made e.g.
-in [4], does not impact the hardness of $\mathcal{E}$-SDP.
-$\mathbf{H}=(\mathbf{A},\mathbf{1}_{r})$, where $\mathbf{A}\in\mathbb{F}^{r\times k}$ and $\mathbf{1}_{r}\in\mathbb{F}^{r\times r}$ denotes the identity matrix.
-Then, $\mathbf{e}\in\mathbb{F}^{n}$ satisfies $\mathbf{e}\mathbf{H}^{\top}=\mathbf{s}$ if and only if it is of the form $\mathbf{e}=(\mathbf{w},\mathbf{s}-\mathbf{w}\mathbf{A}^{\top})$ with $\mathbf{w}\in\mathbb{F}^{k}$.
+We assume that the parity-check matrix is in systematic form  [^5] This assumption does not impact security.
+$\mathbf{H}=(\mathbf{A},\mathbf{1}_{r})$, where $\mathbf{A}\in\mathbb{F}_{p}^{r\times k}$ and $\mathbf{1}_{r}\in\mathbb{F}_{p}^{r\times r}$ denotes the identity matrix.
+Then, $\mathbf{e}\in\mathbb{F}_{p}^{n}$ satisfies $\mathbf{e}\mathbf{H}^{\top}=\mathbf{s}$ if and only if it is of the form $\mathbf{e}=(\mathbf{w},\mathbf{s}-\mathbf{w}\mathbf{A}^{\top})$ with $\mathbf{w}\in\mathbb{F}_{p}^{k}$.
 That is, the witness is defined as a length-$k$ partial error vector that is expanded to a length-$n$ error vector so that the syndrome equations are satisfied.
 […]
-Overall, we obtain the system of polynomial constraints $F(\mathbf{x})=(f_{1},\ldots,f_{n})\in\mathbb{F}[x_{1},\ldots,x_{k}]^{n}$ with
-$f_{i}(\mathbf{x})=f_{\mathcal{E}}\left(x_{i}\right)\quad\text{for }i\in[k],\quad\text{and}\quad f_{k+i}(\mathbf{x})=f_{\mathcal{E}}\left(s_{i}-\langle\mathbf{a}_{i},\mathbf{x}\rangle\right)\quad\text{for }i\in[r],$ (1)
-where $\mathbf{a}_{i}\in\mathbb{F}^{k}$ denotes the $i$-th row of $\mathbf{A}$. The correctness and degree of this modeling are stated in the following proposition.
-Proposition 1
-The polynomial relation $F$ given in Equation 1 provides a degree-$z$ modeling of $\mathcal{E}$-SDP: If $\mathbf{w}\in\mathbb{F}^{k}$ satisfies $F\left(\mathbf{w}\right)=\mathbf{0}$, then the error vector $\mathbf{e}=(\mathbf{w},\mathbf{s}-\mathbf{w}\mathbf{A}^{\top})$ solves the $\mathcal{E}$-SDP instance $(\mathbf{H},\mathbf{s})$ with $\mathbf{H}=(\mathbf{A},\mathbf{1}_{r})$.
-Proof
+Overall, we obtain the system of polynomial constraints $F(\mathbf{x})=(f_{1},\ldots,f_{n})\in\mathbb{F}_{p}[x_{1},\ldots,x_{k}]^{n}$ with
+$f_{i}(\mathbf{x})$ $=f_{\mathcal{E}}\left(x_{i}\right)\quad\text{for }i\in[k],$
+$\text{and}\quad f_{k+i}(\mathbf{x})$ $=f_{\mathcal{E}}\left(s_{i}-\langle\mathbf{a}_{i},\mathbf{x}\rangle\right)\quad\text{for }i\in[r],$ (2)
+where $\mathbf{a}_{i}\in\mathbb{F}_{p}^{k}$ denotes the $i$-th row of $\mathbf{A}$. The correctness and degree of this modeling are stated in the following proposition.
+Proposition 2.
+The polynomial relation $F$ given in Equation˜2 provides a degree-$z$ modeling of $\mathcal{E}$-SDP: $\mathbf{w}\in\mathbb{F}_{p}^{k}$ satisfies $F\left(\mathbf{w}\right)=\mathbf{0}$, if and only if the error vector $\mathbf{e}=(\mathbf{w},\mathbf{s}-\mathbf{w}\mathbf{A}^{\top})$ solves the $\mathcal{E}$-SDP instance $(\mathbf{H},\mathbf{s})$ with $\mathbf{H}=(\mathbf{A},\mathbf{1}_{r})$.
+Proof.
 For $i\in[k]$, $f_{i}(\mathbf{x})$ has degree $z$.
@@
 i.e., the syndrome equations are satisfied. Further, $f_{\mathcal{E}}\left(e_{i}\right)=0$ for all entries of $\mathbf{e}$, i.e., $\mathbf{e}\in\mathcal{E}^{n}$. ∎
-3.0.4 Performance in TCitH framework.
-From the provided modeling of $\mathcal{E}$-SDP, a digital signature is derived using the TCitH framework [21], as summarized in Section 2.0.4.
-The chosen parameters and the obtained signature sizes are provided in Table 1.
+$\mathcal{E}$-SDP-enabled tradeoffs.
+For the proposed modeling, the signature sizes in both TCitH and VOLEitH scale as $\left\lvert\sigma_{F}\right\rvert\propto z$ and $\left\lvert\sigma_{\mathbf{w}}\right\rvert\propto k\cdot\log_{2}(p)$, while $\sigma_{\mathrm{sym}}$, composed of symmetric primitives, is largely independent of the modeling.
+Figure˜2 shows the tradeoff between $k\cdot\log_{2}(p)$ and $z$ which can be achieved by different parameterizations of $\mathcal{E}$-SDP.
+The solver cost is approximated by Proposition˜1.
+This approximation is tight for both Ternary-SDP and CROSS-SDP, parametrized according to the careful analysis in [23] and [21], respectively.
+We observe that only a minor improvement in witness size may be possible by increasing the degree of the relation (i.e., the size of $\mathcal{E}$).
+For fixed $z$, smaller field sizes $p$ are preferable.
+In the following, we focus on CROSS-SDP and Ternary-SDP; by Figure˜2, the latter is expected to yield the smallest $\mathcal{E}$-SDP-based signatures.
+[…]
+Figure 2: Tradeoff between degree and witness size.
+Black lines correspond to $\mathcal{E}$-SDP parameters achieving NIST security category 1, as per Equation˜1.
+Ternary-SDP and CROSS-SDP are parametrized according to [23] and [21].
+TABLE II: Parameters of signatures based on $\mathcal{E}$-SDP obtained via the VOLEitH framework.
+The proof system parameters are selected analogously to those in SDitH [3] to ensure comparable runtime performance of the VC scheme.
+Security category $\mathcal{E}$-SDP Proof System Performance
+[…]
+NIST 1 $127$ 77 $127$ $76$ $16$ $256$ $32$ $101$ 33 fast $6432$
+$11$ $2048$ $22$ $107$ $10$ short $4372$
+33 22 $579$ $213$ $16$ $256$ $96$ $101$ 11 fast $3515$
+$11$ $2048$ $81$ $107$ 88 short $2974$
+NIST 3 $127$ 77 $187$ $111$ $24$ $256$ $48$ $153$ 33 fast 14 35914\,359
+$16$ $4096$ $32$ $157$ 33 short $9361$
+33 22 $839$ $309$ $24$ $256$ $144$ $153$ 11 fast $7816$
+$16$ $4096$ $128$ $157$ 11 short $6463$
+NIST 5 $127$ 77 $251$ $150$ $32$ $256$ $64$ $206$ 33 fast 25 57325\,573
+2121 $4096$ $42$ $216$ 77 short 16 58916\,589
+33 22 $1102$ $406$ $32$ $256$ $192$ $206$ 11 fast 13 85113\,851
+2121 $4096$ $168$ $216$ 55 short 11 52111\,521
+IV Concrete Performance Evaluation
+Performance TCitH.
+From the provided modeling of $\mathcal{E}$-SDP, a digital signature is derived using the TCitH framework [8] and the Fiat-Shamir Transform, as summarized in Section˜II.
+The chosen parameters and the obtained signature sizes are provided in Table˜I.
 The parameters $\tau$, $N$, and $w$ of the proof system are selected following [6] to allow for comparable runtime performance of the VC scheme.
-To simplify implementation, the smallest extension degree $\mu=[\mathbb{K}:\mathbb{F}]$ satisfying the requirement $N\leq\left\lvert\mathbb{K}\right\rvert=\left\lvert\mathbb{F}\right\rvert^{\mu}$ is chosen.
-This, in turn, determines the batching parameter $\eta$.
-Table 1: Parameters of signatures based on $\mathcal{E}$-SDP obtained via the TCitH framework.
-The proof system parameters are selected analogously to those in MQOM [14] to ensure comparable runtime performance of the VC scheme.
-Security
-category
-$\mathcal{E}$-SDP Proof System Performance
-[…]
-NIST 1
-$127$
-77
-$127$
-$76$
-$24$ $256$ 22 $10$ 44 fast $7650$
-$15$ $2048$ 22 $10$ 66 short $5533$
-33
-22
-$579$
-$213$
-$17$ $256$ 66 $14$ 99 fast $3533$
-$12$ $2048$ 77 $12$ 88 short $3095$
-NIST 3
-$127$
-77
-$187$
-$111$
-$36$ $256$ 22 $14$ 66 fast 16 67516\,675
-$23$ $2048$ 22 $14$ 44 short 12 35412\,354
-33
-22
-$839$
-$309$
-$27$ $256$ 66 2121 33 fast $8284$
-$18$ $2048$ 77 $18$ $12$ short $6860$
-NIST 5
-$127$
-77
-$251$
-$150$
-$48$ $256$ 22 1919 77 fast 29 83929\,839
-$31$ $2048$ 22 1919 33 short 22 30522\,305
-33
-22
-$1102$
-$406$
-$36$ $256$ 66 $27$ 44 fast 14 58414\,584
-$25$ $2048$ 77 $24$ 66 short 12 60812\,608
+To simplify implementation, the smallest extension degree $\mu=[\mathbb{K}:\mathbb{F}_{p}]$ satisfying the requirement $N\leq\left\lvert\mathbb{K}\right\rvert=\left\lvert\mathbb{F}_{p}\right\rvert^{\mu}$ is chosen.
+This determines the batching parameter $\eta$.
 For Ternary-SDP, parameterized to achieve NIST security category 1, a signature size as small as $3.1\text{\,}\mathrm{kB}$ is possible.
@@
 Nevertheless, we obtain a significant reduction compared to CROSS [12], for which the smallest signature size is $12.4\text{\,}\mathrm{kB}$.
-3.0.5 Performance in VOLEitH framework.
-Table 2 provides an overview of the performances that can be achieved if TCitH is replaced by the VOLEitH framework.
+Performance VOLEitH.
+Table˜II provides an overview of the performances that can be achieved if TCitH is replaced by the VOLEitH framework.
 Parameters of the proof system are chosen analogously to SDitH [3].
@@
 A similar effect is observed in [6]: for OWFs with small degree and small witness size, the structurally simpler TCitH framework remains competitive.
-On the other hand, for CROSS-SDP, which has a degree of $d=z=7$, VOLEitH enables a reduction in signature size of about $1\text{\,}\mathrm{kB}$ for NIST security category 1, and of up to $6\text{\,}\mathrm{kB}$ for NIST security category 5.
-Table 2: Parameters of signatures based on $\mathcal{E}$-SDP obtained via the VOLEitH framework.
-The proof system parameters are selected analogously to those in SDitH [3] to ensure comparable runtime performance of the VC scheme.
-Security
-category
-$\mathcal{E}$-SDP Proof System Performance
-[…]
-NIST 1
-$127$
-77
-$127$
-$76$
-$16$ $256$ $32$ $101$ 33 fast $6432$
-$11$ $2048$ $22$ $107$ $10$ short $4372$
-33
-22
-$579$
-$213$
-$16$ $256$ $96$ $101$ 11 fast $3515$
-$11$ $2048$ $81$ $107$ 88 short $2974$
-NIST 3
-$127$
-77
-$187$
-$111$
-$24$ $256$ $48$ $153$ 33 fast 14 35914\,359
-$16$ $4096$ $32$ $157$ 33 short $9361$
-33
-22
-$839$
-$309$
-$24$ $256$ $144$ $153$ 11 fast $7816$
-$16$ $4096$ $128$ $157$ 11 short $6463$
-NIST 5
-$127$
-77
-$251$
-$150$
-$32$ $256$ $64$ $206$ 33 fast 25 57325\,573
-2121 $4096$ $42$ $216$ 77 short 16 58916\,589
-33
-22
-$1102$
-$406$
-$32$ $256$ $192$ $206$ 11 fast 13 85113\,851
-2121 $4096$ $168$ $216$ 55 short 11 52111\,521
-3.0.6 Comparison with NIST submissions.
-A comparison of the achieved signature sizes with submissions to NIST’s call for additional digital signature schemes [24] is provided in Table 3.
+On the other hand, for CROSS-SDP, which has a degree of $d=z=7$, VOLEitH enables a reduction in signature size of about $1\text{\,}\mathrm{kB}$ for NIST category 1, and of up to $6\text{\,}\mathrm{kB}$ for NIST category 5.
+Comparison with NIST submissions.
+A comparison of the achieved signature sizes with other ZK-based signatures submissions to NIST’s call for additional digital signature schemes [25] is provided in Table˜III.
 As can be seen, the TCitH and VOLEitH frameworks can be combined with a wide range of OWFs.
 The proposed modeling of $\mathcal{E}$-SDP is competitive compared to the state-of-the-art: for the Ternary-SDP, we obtain a signature size which is very close to those of the smallest proposals that utilize the TCitH or the VOLEitH framework with an alternative hardness assumption.
-The sizes obtained from CROSS-SDP are slightly larger than those of SDitH and PERK, which also use modelings of degree greater than two, but smaller than $z=7$.
-The code-based proposals CROSS and LESS target other trade-offs: CROSS uses a simple CVE-like protocol [18, 6], and LESS achieves a small signature size by using multiple public keys (resulting in a public-key size that is multiple times larger than those of the other schemes shown).
-Table 3: Comparison of the obtained signature sizes with NIST submissions. The shown sizes correspond to the short optimization tradeoff; runtime may vary.
-Scheme Hardness Assumption Design Principle $\left\lvert\sigma\right\rvert$ in $\mathrm{kB}$]
+The sizes obtained from CROSS-SDP are slightly larger than those of SDitH [3] and PERK [4], which also use modelings of degree greater than two, but smaller than $z=7$.
+The code-based proposals CROSS and LESS target other trade-offs: CROSS uses a simple CVE-like protocol [26, 12], and LESS achieves a small signature size by using multiple public keys (resulting in a public-key size that is multiple times larger than those of the other schemes shown) [11].
+SQISign [15] is significantly more compact than the other schemes; this comes at the cost of a complex and slow construction.
+TABLE III: Comparison with other ZK-based signatures. Shown sizes are for the short optimization tradeoff; runtime may vary.
+Scheme Assumption Design $\left\lvert\sigma\right\rvert$ in $\mathrm{kB}$]
 this work CROSS-SDP VOLEitH $4.4$
 Ternary-SDP TCitH $3.1$
+[27] CROSS-$\mathcal{E}$-SDP MPCitH $5.5$
 CROSS CROSS-$\mathcal{E}$-SDP CVE $12.4$
@@
 MQOM Multivariate Quadratic TCitH $2.8$
-FEAST AES VOLEitH $4.5$
+FAEST AES VOLEitH $4.5$
 AES-EM $3.9$
-4 Conclusion
+SQISign Supersingular Isogenies sigma protocol $0.15$
+Comparison with [27].
+Recently [27] proposed a new MPCitH-based protocol for $\mathcal{E}$-SDP.
+Instead of relying on the TCitH and VOLEitH frameworks, their construction is built from the ground up, designing an MPC protocol that computes the syndrome corresponding to a shared restricted vector and turning it into a digital signature using the classical MPCitH framework.
+The approach presented in this paper offers greater flexibility in parameter selection and achieves smaller signatures, in particular through Ternary-SDP.
+V Conclusion
 This paper presents a polynomial modeling of the restricted decoding problem $\mathcal{E}$-SDP, enabling the construction of digital signatures in the TCitH and the VOLEitH frameworks.